OpenStudy (anonymous):

hi , i am having a lot of trouble with this problem any help would be appreciated A little cat, Bella, walks in a straight line, which we shall call the x axis, with the positive direction to the right. As an observant scientist, you make measurements of her motion and construct a graph of the little feline's velocity as a function of time. Find the x component of Bella's velocity at t = 4.00 s. give me a sec have to draw graph...

OpenStudy (anonymous):

to see the graph i found the question online here it is http://www.cobaltss.net/forums/lounge-42/anyone-good-physics-151699/

OpenStudy (anonymous):

\[v(t)=\frac{-4}{3}t+8\]

OpenStudy (anonymous):

\[a(t)=\frac{-4}3{}\]

OpenStudy (anonymous):

a) t=4 v(t)=-16/3+8 v(t=4)=2.66

OpenStudy (anonymous):

can you please explain to me what the question is asking

OpenStudy (anonymous):

i mean how did you figure out the first equation you wrote

OpenStudy (anonymous):

ok, the line is passing (0,8) and (6,0) right

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

and we know its slope is -8/6

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

y=mx+n its formula

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

put m=-8/6 and (0,8) or (6,0) to find n

OpenStudy (anonymous):

y=-8/6x+8=-4/3+8

OpenStudy (anonymous):

y=V(t) x=t

OpenStudy (anonymous):

so this is how find the velocity

OpenStudy (anonymous):